If the area of a rectangle is 1/12 square yds, what is the width if the length is 2yds? Also explain the answer...thanks
Answer by Alan3354(69443)

If the area of a rectangle is 1/12 square yds, what is the width if the length is 2yds? Also explain the answer...thanks
--------------
Area = L*W
1/12 = W*2
W = 1/24 yd